OpenStudy (anonymous):

can u explain 3x+20=80

OpenStudy (bahrom7893):

subtract 20 from both sides: 3x+20-20=80-20

OpenStudy (bahrom7893):

3x + 0 = 60 3x = 60

OpenStudy (bahrom7893):

divide both sides by 3: (3/3)x = 60/3 x = 20

OpenStudy (bahrom7893):

Please click on Become a FAN if I helped, thanks! (Im tryin to get to 100)

OpenStudy (anonymous):

Distribute (or multiply) 2 to 3x and -20, eliminating your parentheses. thus you would have... 2x + 6x - 40 = 80. Then join your like terms together to make your equation shorter. 2x + 6x and you end up with... 8x - 40 = 80 then start to isolate your variable! add 40 to -40 and 80. 8x = 120 now divide 8 to both sides, isolating x! x = 15

OpenStudy (anonymous):

plz tell me if that help at all